OpenStudy (anonymous):

Which answer shows the FOIL method used to expand the product below? (2x+3)(2x2-5)

OpenStudy (anonymous):

F: 2x * 2x^2 = 4x^3 O: 2x * -5 = -10x I: 3 * 2x^2 = 6x^2 L: 3 * -5 = -15 Answer is total of these: 4x^3 - 10x + 6x^2 - 15

OpenStudy (anonymous):

nope thats not one of the answer choices

OpenStudy (anonymous):

did you mean 2x instead of 2x^2 here (2x+3)(>>2x2<<-5)